### v1

# On Creating a Causally Grounded Usable Rating Method for Assessing the Robustness of Foundation Models Supporting Time Series

Abstract
Foundation Models (FMs) have improved time series forecasting in various sectors, such as finance, but their vulnerability to input disturbances can hinder their adoption by stakeholders, such as, investors and analysts. To address this, we propose a causally grounded rating framework to study the robustness of Foundational Models for Time Series (FMTS) with respect to input perturbations. We evaluate our approach on the stock price prediction problem, a well studied problem with easily accessible public data, evaluating six state-of-the-art (some multi-modal) FMTS across six prominent stocks spanning three industries. The ratings proposed by our framework effectively assess the robustness of FMTS and also offer actionable insights for model selection and deployment. Within the scope of our study, we find that (1) multi-modal FMTS exhibit better robustness and accuracy compared to their uni-modal versions and, (2) FMTS pre-trained on time series forecasting task exhibit better robustness and forecasting accuracy compared to general-purpose FMTS pre-trained across diverse settings. Further, to validate our framework’s usability, we conduct a user study showcasing FMTS prediction errors along with our computed ratings. The study confirmed that our ratings reduced the difficulty for users in comparing the robustness of different systems.
1 Introduction
Time series (TS) forecasting uses historical data indexed by time to predict future values. This task finds wide applicability in industry in domains like finance, healthcare, manufacturing, and weather. Although well-studied, the TS forecasting has seen recent advancements with new AI-based approaches including gradient boosting, deep learning, transformers and Foundation Models (FMs) trained on uni-modal numerical data as well as multi-modal data vying for state-of-the-art performance (Elsayed et al. 2021; Jin et al. 2023).
However, having good performance is no guarantee that users will trust a method or model and use it. In particular, users care about the model’s robustness to noisy data and perturbations, as erroneous predictions can have far-reaching impact on stakeholders. The perturbations may have been caused unintentionally by an actor or intentionally by an adversary, but regardless, the users expect robust and consistent performance. To manage user trust, a promising idea is of third-party assessment of models and ratings (automated certifications), which can help users make informed decisions, even without access to the method’s code or model’s training data using both statistical (Srivastava and Rossi 2018, 2020; Srivastava et al. 2024) and causality-based methods (Lakkaraju et al. 2023; Lakkaraju, Srivastava, and Valtorta 2024).
In this context, our contributions are that we: (a) introduce a novel workflow for assessing and rating FMTS for robustness through causal analysis. (b) introduce three perturbations for both numerical and line plots (image) data inspired by real-world applications in unintended scenarios. (c) introduce two novel metrics to measure the causal impact of other attributes on the FMTS (along with existing metrics from literature). (d) create ratings to compare the models in terms of forecasting accuracy and robustness. (e) conduct a user study to assess the ease of interpreting FMTS behavior through our ratings and to assess its alignment with users’ perceptions. With our causally grounded rating framework, we evaluate leading FMTS models—two general-purpose-data trained (Gemini-V and Phi-3 in both uni-modal and multi-modal forms) and two time series-data trained (MOMENT and Chronos) - across diverse architectures (encoder-only, decoder-only, and encoder-decoder) and parameter sizes (46M to 32B), along with three baseline models (ARIMA, random, and biased) (total of 9 models). We use one year of stock price data from six leading companies across three different industries as test data. We find that for both Gemini and Phi-3, their multi-modal versions, in general, exhibit better robustness and forecasting accuracy compared to their uni-modal versions (Figure 6). We also find that time series-specific FMTS exhibit better robustness and forecasting accuracy compared to general-purpose FMTS (Figure 7). Furthermore, the user study confirms that our ratings reduced the difficulty for users in comparing the robustness of different systems.
2 Related Work
Foundation Models Supporting Time Series The use of FMs for time series forecasting has advanced significantly. (Lu et al. 2022) showed that transformers pre-trained on text data can solve sequence modeling tasks in other modalities, enabling their application to time series analysis. Recent studies have reprogrammed LLMs for time series tasks through parameter-efficient fine-tuning and tokenization strategies (Zhou et al. 2023; Gruver et al. 2023; Jin et al. 2023; Cao et al. 2023; Ekambaram et al. 2024). (Zhou et al. 2023) and (Jin et al. 2023) further illustrate the versatility and robustness of fine-tuned language pre-trained transformers for diverse time series tasks. Several models have advanced time series forecasting. (Ansari et al. 2024) and (Woo et al. 2024) have improved forecasting accuracy and model generalization, while (Rasul et al. 2023) and (Das et al. 2023) have explored new tokenization strategies and fine-tuning methods. (Garza and Mergenthaler-Canseco 2023) and (Ekambaram et al. 2024) developed lightweight models for real-time applications, and (Talukder, Yue, and Gkioxari 2024) integrated multiple temporal patterns to improve precision. FMs trained from scratch, like (Gruver et al. 2023), achieved SOTA in zero-shot forecasting, with (Cao et al. 2023) and (Goswami et al. 2024) further improving model performance. Please see Section 5.1 for the FMTS we selected due to their SOTA performance in their respective categories.
Perturbations in Time Series Data TS data is commonly stored in spreadsheets and databases, which are prone to changes due to acts of omission (e.g., negligence, data-entry errors) or commission (e.g., adversarial attacks, sabotage). Omission errors are most common (Panko and Halverson 1996). Tools like Microsoft Excel and Google Sheets are widely used for data collection and analysis, allowing end-user programming (Birch, Lyford-Smith, and Guo 2018). However, over 90% of spreadsheets contain errors due to issues like incorrect formulae, leading to multi-billion dollar losses (Pak-Lok POON and TANG 2024). Adversarial attacks are also increasing in data stores and AI models for tasks like forecasting. (Karim et al. 2019) adapted these concepts to time series, exploring both black-box and white-box attacks. (Oregi et al. 2018) revealed the vulnerability of distance-based classifiers. (Rathore et al. 2020) examined various adversarial attacks on time series classifiers. TSFool (Li et al. 2022) introduced a multi-objective black-box attack to craft imperceptible adversarial time series to fool RNN classifiers.
Rating AI Systems Several works have assessed and rated AI systems for trustworthiness from a third-party perspective without access to training data. (Srivastava and Rossi 2020) proposed a method to rate AI systems for bias, specifically targeting gender bias in machine translators (Srivastava and Rossi 2018), and used visualizations to communicate these ratings (Bernagozzi et al. 2021b). They conducted user studies on trust perception through visualizations (Bernagozzi et al. 2021a), but these lacked causal interpretation. (Lakkaraju, Srivastava, and Valtorta 2024) introduced a causal analysis approach to rate bias in sentiment analysis systems, extending it to assess their impact when used with translators (Lakkaraju et al. 2023). We extend their method to rate MM-TSFM for robustness against perturbations. Causal analysis offers advantages over statistical analysis by determining accountability, aligning with humanistic values, and quantifying the direct influence of various attributes on forecasting accuracy.
3 Problem
3.1 Preliminaries
Time Series Forecasting Let the time series be represented by {xt-n+1, xt-n+2, …., xt, xt+1, …, xt+d}, where each xt-n+i represents a value in time series, where $n$ is called the sliding window size and $d$ is the number of future values the model predicts. Let Xt === {xt-n+1, xt-n+2, …., xt}, and $\hat{Y_{t}}$ $=$ $\hat{x}$t+1, $\hat{x}$t+2, …., $\hat{x}$t+d}, where $\hat{Y_{t}}=f(X_{t}$; $\theta$) for uni-modal FMTS, and in the case of multi-modal FMTS, $X_{t}$ includes both the numerical time series values and the corresponding time series line plots (images). The function $f$ represents pre-trained FMTS with parameter $\theta$ used in a zero-shot manner without task-specific fine-tuning that predicts the values for the next ‘d’ timesteps based on the values at previous $n$ timesteps. Let $Y_{t}$ denote the true values for the next ‘d’ timesteps. Let $S$ be the set of FMTS we want to rate. Let Rt be the residual for the sliding window [t + 1, t + d] and is computed by ($\hat{Y_{t}}$ $Y_{t}$) at each timestep. Our rating method aims to highlight the worst-case scenario for the model. Therefore, we consider the maximum residual, denoted as $R^{max}_{t}$.
Figure 1: Causal model $\mathcal{M}$ for FMTS. The validity of link ‘1’ depends on the data distribution ($P|Z$), while the validity of the links ‘2’ and ‘3’ are tested in our experiments.
Figure 2: Variants of the causal diagram in Figure 1 used to answer different research questions (RQs).
Causal Model The causal model $\mathcal{M}$, is shown in Figure 1. Arrowheads indicate the causal direction from cause to effect. If Sensitive Attribute ($Z$) is a common cause for both Perturbation ($P$) and Residual ($R^{max}_{t}$), it introduces a spurious correlation between $P$ and $R^{max}_{t}$, known as the confounding effect, making $Z$ the confounder. The path from Perturbation to Residual through the confounder is called a backdoor path and is undesirable. Various backdoor adjustment techniques can remove the confounding effect (Xu and Gretton 2022; Fang et al. 2024; Liu et al. 2021). The deconfounded distribution, after adjustment, is represented as $(R^{max}_{t}|do(P))$. The ‘do(.)’ operator in causal inference denotes an intervention to measure the causal effect of $P$ on the $R^{max}_{t}$. Solid red arrows with ‘?’ in Figure 1 denote the causal links tested in our experiments, while the dotted red arrow represents a potential causal link, depending on the distribution $(P|Z)$ across different values of $Z$.
3.2 Problem Formulation
We aim to answer the following research questions (RQs) (with causal diagrams in Fig 2) through our causal analysis when different perturbations denoted by $P$ = {0, 1, 2, 3} (or simply $P0,P1,P2,P3$) are applied to the input given to the set of FMTS $S$:
RQ1: Does $Z$ affect $R^{max}_{t}$, even though $Z$ has no effect on $P$? That is, if perturbations are independent of the sensitive attribute, can the attribute still affect the system outcome, leading to statistical bias (i.e., lack of fairness)? Causal analysis is unnecessary here due to no confounding effect.
RQ2: Does $Z$ affect the relationship between $P$ and $R^{max}_{t}$ when $Z$ has an effect on $P$? That is, if the applied perturbations depend on the value of the sensitive attribute, would the sensitive attribute add a spurious (false) correlation between the perturbation and the outcome of a system leading to confounding bias?
RQ3: Does $P$ affect $R^{max}_{t}$ when $Z$ may have an effect on $R^{max}_{t}$? That is, what is the impact of the perturbation on the outcome of a system when the sensitive attribute may still have an effect on the outcome of a system?
RQ4: Does $P$ affect the accuracy of $S$? That is, do the perturbations affect the performance of the systems’ accuracy? Causal analysis is not required to answer this question as we only need to compute appropriate accuracy metrics to assess how robust a system is against different perturbations.
4 Solution Approach
Our solution approach consists of the following components: (1) Three perturbations that are applied to both numerical time series data and line plots. (2) Metrics, APE and PIE % that assess the performance of FMTS from accuracy and robustness perspectives and aid in assigning ratings. (3) A ‘data to predictions’ and ‘predictions to ratings’ workflow for rating FMTS using both uni-modal (numerical) and multi-modal (line plots (image)) data.
Perturbations: We introduce one syntactic (STP) and two semantic perturbations (SMP) (Figure 3(a)) inspired by real-world applications in unintended scenarios to assess the accuracy and robustness of FMTS. SMPs alter data meaning while preserving context, e.g., a stock’s value might change due to data entry errors or market-specific catalysts. STPs modify the structure of the data while preserving without altering the content. Drop-to-zero (P1) is an SMP inspired by common data entry errors (Ley et al. 2019). Every $n^{th}$ value in the original stock price data is set to zero. Sampling the time series with a sliding window of size $n$ ensures each sample contains a zero. In Value halved perturbation (P2), an SMP, we reduce every $n^{th}$ number in the original stock price data to half of its value. This perturbation simulates periodic adjustments, possibly reflecting events like stock splits or dividend payments. Missing values perturbation (P3), an STP, converts every $n^{th}$ number in the original stock price data to a null value, simulating real-world missing data points in financial datasets due to system incomplete transmissions.
4.1 Evaluation Metrics
In this section, we describe our evaluation metrics for measuring forecasting accuracy and robustness. [^1] Full formulas are provided in the Appendix C.
Forecasting Accuracy Metrics We evaluate forecasting accuracy using three metrics (Makridakis, Spiliotis, and Assimakopoulos 2022): Symmetric mean absolute percentage error (SMAPE) measures average percentage error between actual and predicted values. Mean absolute scaled error (MASE) compares the mean absolute error of forecasts to a naive one-step forecast. Sign Accuracy quantifies how well predicted forecasts align with recent observed values. For SMAPE and MASE, lower values indicate better performance, while higher values are better for Sign Accuracy.
[…]
Average Perturbation Effect (APE) Average Treatment Effect provides the average difference in outcomes between between treated and untreated units (Wang, Nianogo, and Arah 2017). In our context, it computes the difference between perturbed data residuals (P1 through P3) and the unperturbed data residuals (P0), thereby measuring the impact of the perturbation on the outcome. Hence, we refer to this metric as APE. This metric helps us answer RQ3. It is formally defined using the following equation:
$[|E[R^{max}_{t}=j|do(P=i)]-E[R^{max}_{t}=j|do(P=0)]|]$ (1)
Propensity Score Matching - Deconfounding Impact Estimation % (PSM-DIE % or PIE %) In (Lakkaraju, Srivastava, and Valtorta 2024), a linear regression model was used to estimate causal effects, assuming a linear relationship between variables. This method, however, doesn’t capture non-linear relationships or fully eliminate confounding biases. They proposed the DIE % metric for binary treatment values. Our work uses six treatment (perturbation) values, applying Propensity Score Matching (PSM) (Rosenbaum and Rubin 1983) to target confounding effects by matching treatment and control units based on treatment probability, similar to RCTs and independent of outcome variables (Baser 2007). We modify DIE % to introduce PIE % (PSM-DIE%) to answer RQ2. It is defined as:
$\left[||APE_{o}|-|APE_{m}||\right]*100$ (2)
$APE_{o}$ and $APE_{m}$ represent APE computed before and after applying PSM, respectively. $PIE\%$ measures the true impact of $Z$ on the relationship between $P$ and $R^{max}_{t}$.
4.2 Workflow
Our proposed workflow consists of two parts: Data to Predictions and Predictions to Ratings. In the first part, as shown in Figure 3(a) [^2] All the figures in the paper can be found in a higher resolution in the Appendix D., FMTS process the input and predict the next ‘d’ timesteps. The FMTS and baseline models are described in Section 5.1. In the second part, illustrated in Figure 3(b), we adapt the method from (Lakkaraju, Srivastava, and Valtorta 2024) to rate text-based sentiment analysis systems for bias. We extended their approach to handle our more complex multi-modal data with multiple perturbations, beyond the original textual data and binary treatments. The modified metrics, APE and PIE%, defined in Section 4.1, help manage this complexity. The rating algorithms are detailed in Appendix A. These metrics, referred to as raw scores, establish a partial order for determining final system ratings, which vary by rating level, L.
(a) Data to predictions
(b) Predictions to ratings
Figure 3: (a) Black arrows denote the unperturbed and red arrows indicate the perturbed paths. Dashed lines shows the multi-modal path. The perturbed parts of the plots are highlighted in red. (b) Workflow for performing statistical and causal analysis to compute raw scores and assign final ratings to the test systems
5 Experiments and Results
This section introduces the FMTS used in our experiments, baseline models, test data, and evaluation metrics, including two new metrics for perturbations and confounders. We also present the user study design, responses, and findings.
5.1 Experimental Apparatus
FMTS
We used four FMs in a zero-shot setting: TS forecasting FMs (MOMENT and Chronos) and general-purpose (GP) multimodal FMs (Gemini-V and Phi-3) adapted for TS forecasting. We set $n=80$ and $d=20$. Table 1 provides an overview of the FMTS architectural details.
Model Mode Size Purpose & Arch. Inf. Time (sec/sample)
Gemini 1.5 Flash Multi 32B GP-1A, 1B, Decoder 1.6 (1A); 10.2 (1B)
Phi-3-vision Multi 4.2B GP-2A, 2B, Enc-Dec 19.7 (1A); 26.6 (1B)
MOMENT-large Uni 385M TS-1, Encoder 0.315
Chronos-T5-small Uni 46M TS-2, Enc-Dec 0.811
Table 1: Overview of the architectural details of FMTS. *Best guess in the absence of official information.
1. MOMENT ($S_{m}$) (Goswami et al. 2024) is an open-source FM that handles a variety of tasks such as forecasting, classification, anomaly detection, and imputation in zero-shot and few-shot settings, and it can also be fine-tuned if needed. We use MOMENT-large for our experiments, which utilizes T5-Large encoder (Raffel et al. 2020) as the base architecture.
2. Chronos ($S_{c}$) (Ansari et al. 2024) is a pretrained probabilistic time series model. It tokenizes time series values using scaling and quantization, then trains transformer-based language models via cross-entropy loss. Chronos uses the T5 encoder-decoder architecture, specifically Chronos-T5-Small for our experiments.
3. Gemini-V (Gemini 1.5 Flash) ($S_{g}$, $S_{g}^{ni}$(Team et al. 2023) is an FM under the Gemini series. This model is designed to understand inputs that include text and images, enabling it to generate relevant text responses based on a comprehensive analysis of the combined data. This model processes text and images to generate relevant text responses from combined data. We call the ‘numeric_only’ mode, $S_{g}$, and the ‘numeric_and_vision’ mode, ($S_{g}^{ni}$).
4. Phi-3 (Phi-3-vision-128k-instruct) ($S_{p}$, $S_{p}^{ni}$) (Abdin et al. 2024) is a lightweight, SOTA open multimodal model. It is built upon datasets that include synthetic data and filtered publicly available websites, with a focus on very high-quality, reasoning-dense data in both text and vision. We call the ‘numeric_only’ mode, $S_{p}$, and the ‘numeric_and_vision’ mode, ($S_{p}^{ni}$). Below is the prompt template we used for time-series forecasting using Gemini-V and Phi-3 models (we omit the text highlighted in blue for uni-modal forecasting:
You are a time series forecasting model that only outputs the forecasted numerical values. Input: <time series> Given the input time series for the past 80 time steps and the corresponding time series plot, can you forecast the next 20 time steps? Provide a list of 20 numeric values only. Do not provide any discussion.
Baselines
We consider the following baselines:
1. AutoRegressive Integrated Moving Average (ARIMA) ($S_{a}$) is a widely used statistical approach for time series forecasting. It combines three different components: Autoregressive (AR), differencing (I), and moving average (MA) to capture the patterns in the time-series data and predict the next ‘d’ (from section 3) values.
2. Biased system ($S_{b}$) is an extreme baseline biased towards META and GOOG (technology companies), assigning residuals of 0 and 200 respectively, while assigning higher residuals to other companies, representing maximum bias.
3. Random system ($S_{r}$) assigns random price predictions within a company range for contextually meaningful values.
Test Data
We collected daily stock prices from Yahoo! Finance for six companies across different industries: Meta (META) and Google (GOO) in social technology, Pfizer (PFE) and Merck (MRK) in pharmaceuticals, and Wells Fargo (WFC) and Citigroup (C) in financial services. The data spans from March 28, 2023, to April 22, 2024. We used data from March 28, 2023, to March 22, 2024, to predict stock prices for the following month.
5.2 Experimental Evaluation
Figure 4: Studying each metric with respect to impact of company and industry as confounders for all models and all perturbations. Plotted in double logarithmic scale, lower values indicate better robustness. Ratings generated by our method (with $L=3$) are shown on the top of each plot. The complete final order (with ratings) are shown in Table 2 in Appendix E.
In this section, we describe the experimental setup used to address the RQs stated in Section 3, the results obtained, and the conclusions drawn from the results. Figure 2 shows the causal diagrams used to answer the RQs.
Figure 5: Radar plots showing for all systems (a) mean forecasting accuracy with respect to all metrics, (b) forecasting accuracy under P2 (half-valued pertubation) Appendix E: Figs. 18, 19 show all perturbations; (c) mean robustness metrics for FMTS and $S_{a}$, and (d) robustness under P2. Each axis is normalized and inverted if needed so that outer ring implies better performance.
RQ1: Does Sensitive Attribute affect the Residual, even though Sensitive Attribute has no effect on Perturbation? Setup: In this experiment, the causal link from the Sensitive Attribute to Perturbation is absent, as the perturbation to the stock prices does not depend on the corresponding company name or the industry i.e., perturbations are applied uniformly across all the data points. We quantify the statistical bias exhibited by the systems by using WRS described in Section 4.1. We perform two different analyses in this experiment: one to measure the discrepancy shown across different industries (WRSIndustry) and another to measure the discrepancy among both the companies (WRSCompany) within the same industry. Results and conclusion: From Figure 4, most discrepancies can be observed across industries compared to the discrepancies across companies within each industry. When the input data was subjected to perturbation P3 ($WRS_{avg}$ of 5.35333Table 4 in Appendix E shows the RQs along with the average values for all metrics across different systems and perturbations.), the systems exhibited more statistical bias. From Figure 5, $S_{a}$ ($WRS_{avg}$ of 3.96) exhibited the least statistical bias, while $S_{p}$ ($WRS_{avg}$ of 6.28) exhibited the highest statistical bias among the systems evaluated under the perturbations considered. Hence, we conclude that Sensitive Attribute affects the Residual, even though Sensitive Attribute has no effect on Perturbation.
RQ2: Does Confounder affect the relationship between Perturbation and Residual, when Confounder has an effect on Perturbation?
Setup: In this experiment, we use PIE % defined in equation 2 to compare the APE (defined in equation 1) before and after deconfounding using the PSM technique as the presence of the confounder opens a backdoor path from Perturbation to Residual through the Confounder. The causal link from Confounder to Perturbation will be valid only if the perturbation applied depends on the value of the confounder (i.e. the company or the industry the specific data points belong to). To ensure the probability of perturbation assignment varies with respect to the Confounder across three distributions (DI1 through DI3) in the case of Industry and six different distributions in the case of Company (DC1 through DC6), we implement weighted sampling. For each distribution, weights are configured so that perturbation groups P1 through P3 have a twofold higher likelihood of selection compared to P0 for specific values of the confounder. For example, META in DC1, GOOG in DC2, and so on. This strategy highlights significant cases, although other combinations are possible for further exploration. Results and Conclusion: Figure 4 shows that selecting Company as the confounder leads to greater confounding bias in the systems. In Figure 5, $S_{g}^{ni}$ ($PIE_{avg}\%$ of 523.09) exhibited the least confounding bias, while $S_{p}$ ($PIE_{avg}\%$ of 2677.62) exhibited the most. Systems showed more confounding bias under perturbation P3 ($PIE_{avg}\%$ of 1563.94). Therefore, the Confounder affects the relationship between Perturbation and Residual, particularly when the Confounder influences the Perturbation.
RQ3: Does Perturbation affect the Residual when Sensitive Attribute may have an effect on Residual? Setup: The experimental setup in this experiment is same as that for answering RQ2. To compute the APE, we used PSM described in Section 4.1. PSM allows us to effectively determine the effect of Perturbation on the Residual. For instance, if two matched points belong to the same company but only one was perturbed, any difference in their residuals can be directly attributed to the perturbation itself rather than to other confounding factors. This method provides a clear understanding of the true impact of the Perturbation on the Residual. As our rating method aims to bring the worst possible behavior of the systems, we take the MAX(APE) as the raw score that is used to compute the final ratings. Results and Conclusion: It is undesirable to have a higher APE, as it implies that the perturbation applied can have a significant impact on the residuals of different systems. From Figure 4, when Industry was considered as the confounder, it led to a higher APE. As the outcome of $S_{b}$ depended on the Company (and varied from one company to another), the perturbation did not have any effect on the system. Whereas, when Industry was considered as the confounder, the perturbation appeared influential, resulting in a high APE for $S_{b}$. From Figure 5, perturbations had the least impact on $S_{g}^{ni}$ ($APE_{avg}$ of 6.49) and highest impact on $S_{p}$ ($APE_{avg}$ of 28.53). Among all the perturbations, P1 ($APE_{avg}$ of 21.11) was the most disruptive. Hence, Perturbation affects the Residual when Sensitive Attribute may have an effect on Residual.
RQ4: Does Perturbations degrade the accuracy of S? Setup: For this experiment, we do not use any causal model. We compute the three accuracy metrics widely used in for the task of financial time-series forecasting (Makridakis, Spiliotis, and Assimakopoulos 2022) which were summarized in Section 4.1. Results and Conclusion: From Figure 5, $S_{c}$ exhibited the highest amount of forecasting accuracy in terms of SMAPE (average of 0.05) and MASE (average of 4.67), while $S_{a}$ outperformed all other systems in terms of sign accuracy (average of 58.57). $S_{b}$ consistently predicted the correct directional movement of stock prices, exhibiting high sign accuracy as it was designed to adjust residuals based on specific company stock prices. Perturbations P2, P1, and P0 caused the highest decline in SMAPE (average of 0.08), MASE (average of 8.99), and sign accuracy (average of 49.97), respectively. Hence, Perturbations degrade the accuracy of S.
5.3 Overall Performance Comparison
Now, we provide an overall comparison of the different systems across all metrics to highlight key findings about their performance under various perturbations. Figure 5 shows radar plots with mean forecasting accuracy and robustness metrics, including values under perturbation P2.
1. Clear Domination Signals From Figure 5 and the detailed results from Section 5.2, we can draw the following conclusions: $S_{c}$’s Superiority in Forecasting Metrics: $S_{c}$ consistently outperformed other models in terms of forecasting accuracy metrics, specifically SMAPE and MASE. This indicates that $S_{c}$ is highly effective in predicting stock prices with minimal error. General Superiority Over Biased and Random Systems: All models perform better than the biased and random systems in forecasting metrics. This underscores the importance of using well-designed models over naive or biased approaches. Robustness in PIE % and APE Metrics: According to the average scores, the $S_{g}^{ni}$ system demonstrated superior robustness in PIE % and APE metrics. This suggests that $S_{g}^{ni}$ is more resilient to perturbations and confounding biases compared to other systems.
2. Role of Modality The results from Figure 6 indicate that multimodal FMTS, $S_{g}^{ni}$ and $S_{p}^{ni}$, generally perform better in terms of both robustness and accuracy, suggesting that incorporating multiple data modalities (e.g., numerical and image) can improve the system’s ability to make accurate predictions and remain robust against various perturbations.
Figure 6: Effect of the modalities for $S_{g}$ (left) and $S_{p}$ (right).
3. Role of Confounders Our analysis (Figs. 4 and 5: left) shows that using industry as a confounder introduces more bias, with higher PIEscores indicating significant industry-specific effects on the relationship between perturbations and residuals. Inter-industry comparisons also show more discrepancies, as evidenced by WRS scores. Conversely, the impact of company as a confounder varies. For instance, system $S_{b}$ shows minimal effect from company-specific perturbations, suggesting it is well-tailored to company data, while other systems have higher APE scores, indicating a significant impact on residuals due to company-specific factors.
4. Role of Architecture Our evaluation (Fig. 7, left) indicates that the Time Series architecture generally performs better across several metrics, such as achieving the best values in APE_C, PIE_C, SMAPE, MASE, and WRS_I, suggesting that the TS architecture may be more effective for these specific tasks compared to the general purpose architectures. Fig. 7 (right) shows that decoder-only architecture outperforms others in terms of both accuracy and robustness.
Figure 7: Role of architecture in forecasting accuracy and robustness. Performance is averaged across models within each category. See Table 1.
Overall, our comparison highlights $S_{c}$’s forecasting accuracy, the robustness of multimodal systems against perturbations and confounding biases, and the superiority of well-designed models over naive approaches.
5.4 User Study
We conducted a user study to evaluate the ratings generated by our approach for comparing the behavior of various FMTS based on two key metrics: robustness and statistical fairness (defined as lack of statistical bias). To simplify the evaluation for participants, we converted the generated ratings into rankings (i.e., the system with the highest robustness ranking is the most robust system). The main objective of this study was to validate the following hypotheses:
HP1: Rankings generated by our approach decrease the difficulty of comparing system robustness.
HP2: Rankings generated by our approach decrease the difficulty of comparing system fairness (lack of statistical bias).
HP3: Rankings generated by our method align with users’ rankings for both fairness and robustness.
This IRB-approved study [^4] Institutional details anonymized for reviewing. involved participants evaluating FMTS models forecasting stock prices for companies in various industries, participants were introduced to key concepts including robustness, fairness, and error metrics (maximum residual, mean, and standard deviation of errors) to ensure informed evaluations. All inputs were sought on Likert 5-value scale. The details and complete set of questions are provided in the Appendixes  FLABEL:sec:appendix-study-q.
The study was structured into four panels: a self-assessment on knowledge about time series and financial data, a fairness panel, where fairness was the evaluated metric, and two robustness panels, where robustness was assessed under two different perturbations (P1 and P2). In the fairness panel, participants were presented with graphs depicting the residual values of six different systems and an ideal system using stock price data from the Technology and Pharmaceuticals industries. They were provided with the mean and standard deviation of errors and asked to rank the systems from least to most fair. Participants then rated the difficulty of this ranking task (1 being the most difficult). Subsequently, they were shown the rankings generated by our approach and asked to rate the accuracy of these rankings (1 being the least accurate). Finally, participants were asked to rank the difficulty of comparing the behavior of different systems using our rankings (1 being the most difficult). For the robustness panels, similar questions were posed, with users evaluating systems based on their robustness to different perturbations (P1 and P2). A total of 26 users from academia and industry participated over two weeks. We performed different types of statistical tests to draw significant conclusions; see details in Table 5, 6 in Appendix F. We now discuss the key findings from the tests.
To evaluate HP1, we conducted a paired t-test to compare user responses on difficulty of ranking various systems before and after presenting our rankings. The same participants assessed the difficulty using both the graph representation of fairness and our ranking representation, making the paired t-test appropriate. Paired t-test also accounts for the inherent correlation between the paired rankings, making it suitable to account for potential different perceptions across the two representations. Paired t-tests for each robustness panel indicated a significant difference before (P1:$\mu$ = 2.70, $\sigma$ = 1.06; P2: $\mu$ = 2.65, $\sigma$ = 1.17) and after (P1: $\mu$ = 3.23, $\sigma$ = 1.42; P2: $\mu$ = 3.07, $\sigma$ = 1.44) our rankings were presented with P1:t(26) = -1.89, p = 0.030, and P2: t(26) = -1.62, p = 0.059. Since the p-values $\prec$ 0.1, we confirm HP1 that the ranking generated by our approach significantly reduced the perceived difficulty of comparing different systems. Same approach is used to evaluate HP2. The paired t-test showed no significant change in perceived difficulty scores before ($\mu$ = 2.54, $\sigma$ = 1.30) and after ($\mu$ = 2.92, $\sigma$ = 1.35) our rankings were presented, t(26) = -1.18, p = 0.12. Since p $\succ$ 0.1, we conclude that our ranking representation did not significantly reduce the perceived difficulty of comparing different systems. The lack of significant reduction in perceived difficult may have stemmed from the complexity of the graphical representations.
To validate HP3, we used the Spearman Rank Correlation coefficient (Zar 2005) ($\rho$) to evaluate the alignment between the users’ rankings and those produced by our approach. We considered a confidence interval of 90 %. The fairness panel showed a high correlation ($\rho$= 0.73), and the robustness under P1 showed a strong correlation ($\rho$ = 0.91). However, robustness under P2 showed a weak correlation ($\rho$ = 0.14).
In summary, the results of the user study indicate that the rankings generated by our approach can significantly reduce the difficulty of comparing the robustness of different FMTS systems. However, when the comparison metric is fairness, this reduction is not significant. Additionally, while user rankings align well with our method generated rankings for fairness and robustness under P1, they show a weak correlation for robustness under P2, indicating differing perceptions of the ’value halved’ perturbation. P1 (drop-to-zero) involves a significant semantic change that is easier to spot, whereas P2 (value halved) is also a semantic perturbation but subtler, making it potentially harder to identify. Our current study is preliminary and promising; an avenue for future work is to conduct it at a larger scale.
6 Discussion and Conclusion
Our paper aimed to measure the impact of perturbations and confounders on the outcome of FMTS using stock prices across leading companies and industries. We studied Industry and Company as confounders, motivated by the intuition that stakeholders rely on learning-based systems for stock purchase decisions and would be interested in knowing if model errors depend on stock price ranges. For example, does a model commit more errors predicting META’s stock prices compared to MRK’s? To minimize volatility effects, we performed both intra-industry and inter-industry analyses. In future, we plan to study confounders like seasonal trends and financial news. As demonstrated, we believe metrics should be selected based on the questions one wants to answer, rather than relying solely on statistical accuracy. The hypothesis testing approach from (Lakkaraju, Srivastava, and Valtorta 2024), adapted for our work, helped quantify biases and perturbation impacts on test systems. The perturbations used in our analysis have real-world impacts, applicable to both numeric and multi-modal data. While methods like differential evaluation can find the most impactful perturbation variations, we focused on assessing whether simple, subtle perturbations affect FMTS.
Conclusion We proposed a causally grounded empirical framework to study FMTS robustness against three input perturbations, evaluating seven state-of-the-art FMTS across six prominent stocks in three industries. Our framework’s ratings accurately assessed FMTS robustness and provided actionable insights for model selection and deployment. Experiments showed multi-modal FMTS exhibited greater robustness, while uni-modal FMTS had higher forecasting accuracy. FMTS trained on time series tasks showed better robustness and accuracy compared to general-purpose FMTS. A user study confirmed our ratings effectively convey FMTS robustness to end-users, demonstrating the framework’s real-world applicability.
Appendix A Details of Rating Algorithms
Apart from Algorithm 2, the rest were adapted from (Lakkaraju, Srivastava, and Valtorta 2024) to suit the FMTS forecasting setting. Here is how the rating method works.
  1. 1.
    Algorithm 1 computes the weighted rejection score (WRS) which was defined in Appendix C in the main paper.
  2. 2.
    Algorithm 2 computes the PIE % based on Propensity Score Matching (PSM) which was defined in Section 4.1 in the main paper.
  3. 3.
    Algorithm 3 creates a partial order of systems within each perturbation based on the raw scores computed. It will arrange the systems in ascending order w.r.t the raw score. The final partial order (PO) will be a dictionary of dictionaries.
  4. 4.
    Algorithm 4 computes the final ratings for systems within each perturbation based on the PO from previous algorithm. It splits the set of raw score values obtained within each perturbation into ‘L’ parts where ‘L’ is the rating level chosen by the user. Each of the systems is given a rating based on the compartment number in which its raw score belongs. The algorithm will return a dictionary with perturbations as keys and ratings provided to each system within the perturbation as the value.
Purpose: is used to calculate the weighted sum of the number of rejections of null-hypothesis for Dataset $d_{j}$ pertaining to a system $s$, Confidence Intervals (CI) $ci_{k}$ and Weights $w_{k}$.
Input:
$d$, dataset corresponding to a specific perturbation.
$CI$, confidence intervals (95%, 70%, 60%).
$s$, system for which WRS is being computed.
$W$, weights corresponding to different CIs (1, 0.8, 0.6).
Output:
$Z$, Sensitive attribute
$\psi$, weighted rejection score. $\psi\leftarrow 0$ for each $ci_{i},w_{i}\in CI,W$ do
       // $z_{m},z_{n}$ are classes of $Z$
      for each $z_{m},z_{n}\in Z$ do
             $t,pval,dof\leftarrow T-Test(z_{m},z_{n})$;
             $t_{crit}\leftarrow LookUp(ci_{i},dof)$;
             if $t_{crit}>t$ then
                   $\psi\leftarrow\psi+0$;
            else
                  $\psi\leftarrow\psi+w_{i}$
             end if
       end for
end for
return $\psi$
Algorithm 1 WeightedRejectionScore
Purpose: is used to calculate the Deconfounding Impact Estimation using Propensity Score Matching (PIE).
Input:
$s$, a system belonging to the set of test systems, $S$.
$D$, datasets pertaining to a perturbation (different distributions).
$p$, A perturbation other than $p_{0}$
$p_{0}$, control perturbation (or no perturbation.
Output:
$\psi$, PIE score.
$\psi\leftarrow 0$
$PIE\_list\leftarrow[]$             // To store the list of PIE % of all the datasets.
for each $d_{j}\in D$ do
       $APE\_o\leftarrow E(R|P=p_{m})-E(R|P=p_{0})$;
       $APE\_m\leftarrow E(R|do(P=p_{m}))-E(R|do(P=p_{0}))$;
       $PIE\_list[j]\leftarrow(APE\_m-APE\_o)*100$;
end for
$\psi\leftarrow MAX(PIE\_list)$;
return $\psi$
Algorithm 2 ComputePIEScore
Purpose: is used to create a partial order based on the computed weighted rejection score / the PIE %.
Input:
$S$, Set of systems.
$P$, Set of perturbations.
$F$, Flag that says whether the confounder is present (1) or not (0).
$D$, $CI$, $W$ (as defined in the previous algorithms).
Output:
$PO$, dictionary with a partial order for each perturbation.
$PO\leftarrow\{\}$;
$SD\leftarrow\{\}$;
if F == 0 then
       for each $p_{i}\in P$ do
             for each $s_{j}\in S$ do
                   $\psi\leftarrow WeightedRejectionScore(p_{i},s_{j},D,CI,W)$;
                   $SD[s_{j}]\leftarrow\psi$;
             end for
            $PO[p_{i}]\leftarrow SORT(SD)$;
       end for
else
      for each $p_{j}\in P$ do
             for each $s_{i}\in S$ do
                   $\psi\leftarrow ComputePIEScore(s_{i},p_{j},p_{0},D,CI,W)$;
                   $SD[s_{j}]\leftarrow\psi$;
             end for
            $PO[p_{i}]\leftarrow SORT(SD)$;
       end for
end if
return $PO$
Algorithm 3 CreatePartialOrder
Purpose: AssignRating assigns a rating to each of the SASs based on the partial order and the number of rating levels, $L$.
Input:
$S$, $D$, $CI$, $W$, $P$ (as defined in the previous algorithms).
$L$, rating levels chosen by the user.
Output:
$R$, dictionary with perturbations as keys and ratings assigned to each system within each perturbation as the values.
$R\leftarrow\{\}$;
$PO\leftarrow\text{CreatePartialOrder}(S,D,CI,W,G)$;
for $p_{i}\in P$ do
[…]
      if $\text{len}(S)>1$ then
             $G\leftarrow\text{ArraySplit}(\psi,L)$;
             for $k,i\in PO[p_{i}],\psi$ do
                   for $g_{j}\in G$ do
                         if $i\in g_{j}$ then
                               $SD[k]\leftarrow j$;
                         end if
                   end for
             end for
      else
             // Case of a single SAS in $S$
             if $\psi==0$ then
                   $SD[k]\leftarrow 1$;
            else
                   $SD[k]\leftarrow L$;
             end if
       end if
      $R[p_{i}]\leftarrow SD$;
end for
return $R$;
Algorithm 4 AssignRating
Appendix B Additional Related Work
B.1 Robustness Testing of Foundation Models
(Zhang and Ré 2022) examined group distribution shifts and evaluated FMs on image classification tasks with spurious confounders. In our work, we assess the robustness of FMs within time series forecasting by measuring their performance in the presence of two confounders across various perturbation settings and test dataset distributions. (Zhang et al. 2023) used foundation models as a surrogate oracle to measure the robustness of image classification models. However, their test systems did not include any foundation models. (Xiao et al. 2024) introduces a framework, RITFIS, to assess the LLM-based software against natural language input. However, they did not consider any other modalities. (Schlarmann and Hein 2023) showed that imperceivable attacks on images to change the caption output of multi-modal FMs can lead to broadcasting of fake information to honest users. They only evaluate robustness of OpenFlamingo model under different attacks but does not compare its performance with any other FMs. None of these works assess the effectiveness and usability of their robustness testing methods. We address this gap through a user study.
Appendix C Evaluation Metrics
In this section, we define our evaluation metrics: forecasting accuracy and robustness.
Forecasting Accuracy Metrics
We evaluate the systems’ forecasting accuracy using established metrics commonly applied in time-series forecasting tasks (Makridakis, Spiliotis, and Assimakopoulos 2022).
Symmetric mean absolute percentage error (SMAPE) is defined as,
$SMAPE=\frac{1}{T}\sum_{t=1}^{T}\frac{|x_{t}-\hat{x}_{t}|}{(|x_{t}|+|\hat{x}_{t}|)/2},$ (3)
where $T=20$ (i.e., the value of $d$) is the total number of observations in the predicted time series. SMAPE scores range from 0 to 2, with lower scores indicating more precise forecasts.
Mean absolute scaled error (MASE) measures the mean absolute error of forecasts relative to that of a naive one-step forecast on the training data.
$MASE=\frac{\frac{1}{T}\sum_{i=t+1}^{t+T}|x_{i}-\hat{x_{i}}|}{\frac{1}{t}\sum_{i=1}^{t}|x_{i}-x_{i-1}|},$ (4)
where in our case, $t=80$, and $T=100$. Lower MASE values indicate better forecasts.
Sign Accuracy quantifies the average classification accuracy across all test samples, where a higher accuracy indicates more precise predictions. This metric classifies based on how the predicted forecasts align with the most recent observed values in the input time series.
Robustness Metrics We adapt WRS metric originally proposed in (Lakkaraju, Srivastava, and Valtorta 2024) to answer RQ1. Additionally, we introduce two new metrics: APE and PIE % (modified versions of ATE (Abdia et al. 2017) and DIE % (Lakkaraju, Srivastava, and Valtorta 2024)) tailored to answering RQ2 and RQ3.
Weighted Rejection Score (WRS): WRS, introduced in (Lakkaraju, Srivastava, and Valtorta 2024), measures statistical bias. First, Student’s t-test (Student 1908) compares max residual distributions $(R^{max}_{t}|Z)$ for different values of the protected attribute $Z$. We measure this between each pair of industries or companies, resulting in ${}^{3}C_{2}=3$ comparisons. (Lakkaraju, Srivastava, and Valtorta 2024) chose different confidence intervals (CI) [95%, 75%, 60%] that have different critical values and if the computed t-value is less than the critical value, the null hypothesis is rejected. WRS is mathematically defined by the following equation:
[…]
where, $x_{i}$ is the variable set based on whether the null hypothesis is accepted (0) or rejected (1). $w_{i}$ is the weight that is multiplied by $x_{i}$ based on the CI. For example, if CI is 95%, $x_{1}$ is multiplied by 1. The lower the CI, the lower the weight will be. WRS helps us answer RQ1 (see Section 3).
Appendix D Figures in a Higher Resolution from Main Paper
Figure 8: Causal model $\mathcal{M}$ for FMTS. The validity of link ‘1’ depends on the data distribution ($P|Z$), while the validity of the links ‘2’ and ‘3’ are tested in our experiments.
Figure 9: Variants of the causal diagram in Figure 1 used to answer different research questions (RQs).
Figure 10: Data to predictions. Workflow for performing statistical and causal analysis to compute raw scores and assign final ratings to the test systems
Figure 11: Predictions to ratings. Black arrows denote the unperturbed and red arrows indicate the perturbed paths. Dashed lines shows the multi-modal path. The perturbed parts of the plots are highlighted in red.
Figure 12: Studying each metric with respect to impact of company and industry as confounders for all models and all perturbations. Plotted in double logarithmic scale, lower values indicate better robustness. Ratings generated by our method (with $L=3$) are shown on the top of each plot. The complete final order (with ratings) are shown in Table 2 in Appendix E.
Figure 13: Radar plots showing for all systems (a) mean forecasting accuracy with respect to all metrics, (b) forecasting accuracy under P2 (half-valued pertubation) Appendix E: Figs. 18, 19 show all perturbations; (c) mean robustness metrics for FMTS and $S_{a}$, and (d) robustness under P2. Each axis is normalized and inverted if needed so that outer ring implies better performance.
Figure 14: Effect of the modalities for $S_{g}$ (left) and $S_{p}$ (right).
Figure 15: Role of architecture in forecasting accuracy and robustness. Performance is averaged across models within each category. See Table 1.
Appendix E Additional Experimental Results
In this section, Table 2 shows the partial order and final order with respect to all the metrics defined in Section 4.1. Table 3 shows the forecasting accuracy values for all the systems. Table 4 shows the research questions, and average values for all metrics across systems and perturbations (average values are referred to in the conclusions made for each RQ in Section 5.2). Figure 16 and 17 shows bar plots with all the robustness metric values and forecasting accuracy values. Figure 20 shows the heatmap for all metrics for all the models.
Table 2 shows the partial order and final order with respect to all the metrics defined in Section 4.1. Table 3 shows the forecasting accuracy values for all the systems. Table 4 shows the research questions, and average values for all metrics across systems and perturbations (average values are referred to in the conclusions made for each RQ in Section 5.2). Figure 16 and 17 shows bar plots with all the robustness metric values and forecasting accuracy values. Figure 20 shows the heatmap for all metrics for all the models.
Forecasting Evaluation Dimensions P Partial Order Complete Order
[…]
P1 {$S_{a}$: 2.6, $S_{m}$: 4.6, $S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{r}$: 4.6, $S_{p}^{ni}$: 5.9, $S_{p}$: 6.9, $S_{c}$: 6.9, $S_{b}$: 6.9} {$S_{a}$: 1, $S_{m}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{r}$: 1, $S_{p}^{ni}$: 2, $S_{p}$: 3, $S_{c}$: 3, $S_{b}$: 3}
P2 {$S_{a}$: 4.6, $S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{p}^{ni}$: 4.6, $S_{m}$: 4.6, $S_{r}$: 4.6, $S_{c}$: 6.9, $S_{p}$: 6.9, $S_{b}$: 6.9} {$S_{a}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 1, $S_{m}$: 1, $S_{r}$: 1, $S_{c}$: 2, $S_{p}$: 2, $S_{b}$: 2}
P3 {$S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, Sm: 4.6, $S_{r}$: 4.6, $S_{c}$: 4.6, $S_{a}$: 5.9, $S_{p}^{ni}$: 6.9, $S_{p}$: 6.9, $S_{b}$: 6.9} {$S_{g}$: 1, $S_{g}^{ni}$: 1, Sm: 1, $S_{r}$: 1, $S_{c}$: 1, $S_{a}$: 2, $S_{p}$: 3, $S_{p}^{ni}$: 3, $S_{b}$: 3}
[…]
P1 {$S_{a}$: 0.6, $S_{c}$: 4.6, $S_{p}$: 5.9, $S_{p}^{ni}$: 5.9, $S_{r}$: 5.9, $S_{g}^{ni}$: 5.9, $S_{g}$: 6.9, $S_{m}$: 6.9, $S_{b}$: 6.9} {$S_{a}$: 1, $S_{c}$: 1, $S_{p}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 2, $S_{g}^{ni}$: 2, $S_{g}$: 3, $S_{m}$: 3, $S_{b}$: 3}
P2 {$S_{a}$: 2.6, $S_{c}$: 4.6, $S_{r}$: 4.6, $S_{p}^{ni}$: 4.6, $S_{p}$: 5.2, $S_{g}$: 5.9, $S_{g}^{ni}$: 5.9, $S_{m}$: 6.9, $S_{b}$: 6.9} {$S_{a}$: 1, $S_{c}$: 1, $S_{r}$: 1, $S_{p}^{ni}$: 1, $S_{p}$: 2, $S_{g}$: 2, $S_{g}^{ni}$: 2, $S_{m}$: 3, $S_{b}$: 3}
P3 {$S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{p}^{ni}$: 4.6, Sp: 4.6, $S_{c}$: 4.6, $S_{m}$: 6.9, $S_{a}$: 6.9, $S_{r}$: 6.9, $S_{b}$: 6.9} {$S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 1, Sp: 1, $S_{c}$: 1, $S_{m}$: 2, $S_{a}$: 2, $S_{r}$: 2, $S_{b}$: 2}
[…]
P2 {$S_{c}$: 357.72, $S_{g}^{ni}$: 527.76, $S_{g}$: 597.54, $S_{a}$: 1154.87, $S_{m}$: 1326.20, $S_{r}$: 1463.71, $S_{p}^{ni}$: 1653.53, $S_{b}$: 2174.39, $S_{p}$: 2295.68} {$S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{a}$: 2, $S_{m}$: 2, $S_{r}$: 2, $S_{p}^{ni}$: 3, $S_{b}$: 3, $S_{p}$: 3}
P3 {$S_{g}$: 703.94, $S_{g}^{ni}$: 884.34, $S_{c}$: 911.53, $S_{p}^{ni}$: 972.95, $S_{a}$: 1195.04, $S_{m}$: 2998.25, $S_{p}$: 3208.04, $S_{r}$: 3560.94, $S_{b}$: 7489.48} {$S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{c}$: 1, $S_{p}^{ni}$: 2, $S_{a}$: 2, $S_{m}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3}
[…]
P2 {$S_{g}^{ni}$: 469.16, $S_{c}$: 576.18, $S_{g}$: 651.07, $S_{m}$: 1150.45, $S_{a}$: 1275.04, $S_{p}^{ni}$: 2238.21, $S_{p}$: 3257.35, $S_{r}$: 4274.38, $S_{b}$: 9474.61} {$S_{g}^{ni}$: 1, $S_{c}$: 1, $S_{g}$: 1, $S_{m}$: 2, $S_{a}$: 2, $S_{p}^{ni}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3}
P3 {$S_{g}^{ni}$: 436.33, $S_{g}$: 513.47, $S_{c}$: 650.20, $S_{m}$: 866.61, $S_{r}$: 1305.78, $S_{a}$: 1716.68, $S_{b}$: 1846.56, $S_{p}^{ni}$: 2773.74, $S_{p}$: 4064.03} {$S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{c}$: 1, $S_{m}$: 2, $S_{r}$: 2, $S_{a}$: 2, $S_{b}$: 3, $S_{p}^{ni}$: 3, $S_{p}$: 3}
[…]
P2 {$S_{g}$: 3.72, $S_{c}$: 4.79, $S_{g}^{ni}$: 6.06, $S_{a}$: 11.32, $S_{m}$: 13.36, $S_{p}^{ni}$: 18.94, $S_{p}$: 26.02, $S_{r}$: 42.91, $S_{b}$: 101.20} {$S_{g}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{a}$: 2, $S_{m}$: 2, $S_{p}^{ni}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3}
P3 {$S_{a}$: 7.87, $S_{g}$: 8.40, $S_{g}^{ni}$: 9.09, $S_{c}$: 9.50, $S_{p}^{ni}$: 16.73, $S_{m}$: 31.36, $S_{r}$: 36.59, $S_{p}$: 37.39, $S_{b}$: 99.72} {$S_{a}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{c}$: 2, $S_{p}^{ni}$: 2, $S_{m}$: 2, $S_{r}$: 3, $S_{p}$: 3, $S_{b}$: 3}
[…]
P2 {$S_{b}$: 0, $S_{g}^{ni}$: 5.31, $S_{c}$: 6.42, $S_{g}$: 8.69, $S_{p}$: 10.81, $S_{m}$: 13.92, $S_{r}$: 17.61, $S_{a}$: 21.39, $S_{p}^{ni}$: 27.63} {$S_{b}$: 1, $S_{g}^{ni}$: 1, $S_{c}$: 1, $S_{g}$: 2, $S_{p}$: 2, $S_{m}$: 2, $S_{r}$: 3, $S_{a}$: 3, $S_{p}^{ni}$: 3}
P3 {$S_{b}$: 0, $S_{g}^{ni}$: 6.48, $S_{g}$: 7.06, $S_{a}$: 7.42, $S_{c}$: 8.80, $S_{m}$: 10.87, $S_{r}$: 16.63, $S_{p}^{ni}$: 35.35, $S_{p}$: 46.50 } {$S_{b}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{a}$: 2, $S_{c}$: 2,$S_{m}$: 2, $S_{r}$: 3, $S_{p}^{ni}$: 3, $S_{p}$: 3 }
[…]
P1 {$S_{c}$: 0.065, $S_{g}^{ni}$: 0.067, $S_{g}$: 0.072, $S_{a}$: 0.084, $S_{m}$: 0.100, $S_{p}$: 0.100, $S_{p}^{ni}$: 0.100, $S_{r}$: 0.830, $S_{b}$: 1.276 } {$S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{a}$: 2, $S_{m}$: 2, $S_{p}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 3, $S_{b}$: 3 }
P2 {$S_{g}$: 0.051, $S_{c}$: 0.053, $S_{g}^{ni}$: 0.060, $S_{a}$: 0.069, $S_{p}^{ni}$: 0.095, $S_{m}$: 0.098, $S_{p}$: 0.100, $S_{r}$: 0.830, $S_{b}$: 1.276 } {$S_{g}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{a}$: 2, $S_{p}^{ni}$: 2, $S_{m}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
P3 {$S_{a}$: 0.040, $S_{c}$: 0.043, $S_{g}$: 0.049, $S_{g}^{ni}$: 0.056, $S_{p}^{ni}$: 0.078, $S_{p}$: 0.092, $S_{m}$: 0.097, $S_{r}$: 0.830, $S_{b}$: 1.276 } {$S_{a}$: 1, $S_{c}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 2, $S_{p}^{ni}$: 2, $S_{p}$: 2, $S_{m}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
[…]
P1 {$S_{c}$: 5.40, $S_{g}^{ni}$: 5.65, $S_{g}$: 6.13, $S_{p}^{ni}$: 8.87, $S_{p}$: 9.19, $S_{m}$: 9.32, $S_{a}$: 18.36, $S_{r}$: 86.99, $S_{b}$: 947.56 } {$S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{p}^{ni}$: 2, $S_{p}$: 2, $S_{m}$: 2, $S_{a}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
P2 {$S_{g}$: 4.74, $S_{c}$: 4.99, $S_{g}^{ni}$: 5.59, $S_{a}$: 8.24, $S_{p}^{ni}$: 8.49, $S_{m}$: 9.15, $S_{p}$: 9.32, $S_{r}$: 86.87, $S_{b}$: 947.56 } {$S_{g}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{a}$: 2, $S_{p}^{ni}$: 2, $S_{m}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
P3 {$S_{a}$: 3.79, $S_{c}$: 4.10, $S_{g}$: 4.64, $S_{g}^{ni}$: 5.39, $S_{p}^{ni}$: 7.11, $S_{p}$: 8.68, $S_{m}$: 9.03, $S_{r}$: 86.65, $S_{b}$: 947.56 } {$S_{a}$: 1, $S_{c}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 2, $S_{p}^{ni}$: 2, $S_{p}$: 2, $S_{m}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
[…]
P1 {$S_{m}$: 41.19, $S_{p}$: 44.33, $S_{p}^{ni}$: 46.77, $S_{r}$: 49.62, $S_{g}$: 50.53, $S_{c}$: 52.09, $S_{g}^{ni}$: 53.93, $S_{a}$: 57.08, $S_{b}$: 62.60 } {$S_{m}$: 1, $S_{p}$: 1, $S_{p}^{ni}$: 1, $S_{r}$: 2, $S_{g}$: 2, $S_{c}$: 2, $S_{g}^{ni}$: 3, $S_{a}$: 3, $S_{b}$: 3 }
P2 {$S_{m}$: 41.05, $S_{p}$: 44.02, $S_{p}^{ni}$: 47.67, $S_{r}$: 49.64, $S_{g}$: 49.75, $S_{c}$: 50.79, $S_{g}^{ni}$: 54.43, $S_{a}$: 57.13, $S_{b}$: 62.60 } {$S_{m}$: 1, $S_{p}$: 1, $S_{p}^{ni}$: 1, $S_{r}$: 2, $S_{g}$: 2, $S_{c}$: 2, $S_{g}^{ni}$: 3, $S_{a}$: 3, $S_{b}$: 3 }
P3 {$S_{m}$: 40.72, $S_{p}$: 44.26, $S_{p}^{ni}$: 47.50, $S_{r}$: 49.71, $S_{g}$: 51.34, $S_{c}$: 51.35, $S_{g}^{ni}$: 52.97, $S_{a}$: 59.98, $S_{b}$: 62.60 } {$S_{m}$: 1, $S_{p}$: 1, $S_{p}^{ni}$: 1, $S_{r}$: 2, $S_{g}$: 2, $S_{c}$: 2, $S_{g}^{ni}$: 3, $S_{a}$: 3, $S_{b}$: 3 }
Table 2: Final raw scores and ratings based on different metrics computed. Higher ratings indicate higher bias for WRS and PIE %, higher disruption for APE, greater inaccuracy for MASE and SMAPE, and higher accuracy for Sign Accuracy. For simplicity, we denoted the raw scores for accuracy metrics using just the mean value, but standard deviation was also considered for rating. The chosen rating level, L = 3. Overall, across all the settings, system $S_{p}$ exhibited statistical bias in 50 % of cases, confounding bias in 100 % of cases, and disruptive behavior in 50 % of the cases based on APE values.
Metric P $S_{g}$ $S_{g}^{ni}$ $S_{p}$ $S_{p}^{ni}$ $S_{m}$ $S_{c}$ $S_{a}$ $S_{b}$ $S_{r}$
[…]
P1 0.072 $\pm$ 0.123 0.067 $\pm$ 0.178 0.100 $\pm$ 0.125 0.100 $\pm$ 0.143 0.100 $\pm$ 0.076 0.065 $\pm$ 0.189 0.084 $\pm$ 0.282 0.830 $\pm$ 0.639
P2 0.051 $\pm$ 0.047 0.060 $\pm$ 0.085 0.100 $\pm$ 0.088 0.095 $\pm$ 0.097 0.098 $\pm$ 0.074 0.053 $\pm$ 0.092 0.069 $\pm$ 0.217 0.830 $\pm$ 0.639
P3 0.049 $\pm$ 0.045 0.056 $\pm$ 0.052 0.092 $\pm$ 0.074 0.078 $\pm$ 0.078 0.097 $\pm$ 0.071 0.043 $\pm$ 0.048 0.040 $\pm$ 0.037 0.830 $\pm$ 0.640
[…]
P1 6.13 $\pm$ 8.31 5.65 $\pm$ 10.23 9.19 $\pm$ 8.61 8.87 $\pm$ 8.94 9.32 $\pm$ 7.39 5.40 $\pm$ 12.45 18.36 $\pm$ 168.82 86.99 $\pm$ 73.53
P2 4.74 $\pm$ 4.53 5.59 $\pm$ 8.19 9.32 $\pm$ 7.94 8.49 $\pm$ 7.94 9.15 $\pm$ 7.15 4.99 $\pm$ 9.90 8.24 $\pm$ 48.58 86.87 $\pm$ 73.32
P3 4.64 $\pm$ 4.42 5.39 $\pm$ 5.27 8.68 $\pm$ 7.18 7.11 $\pm$ 6.75 9.03 $\pm$ 6.90 4.10 $\pm$ 6.33 3.79 $\pm$ 3.57 86.65 $\pm$ 73.11
Sign Accuracy (%$\uparrow$ P0 52.08 50.41 45.09 47.67 40.70 53.75 60.08 62.60 49.88
P1 50.53 53.93 44.33 46.77 41.19 52.09 57.08 49.62
P2 49.75 54.43 44.02 47.67 41.05 50.79 57.13 49.64
P3 51.34 52.97 44.26 47.50 40.72 51.35 59.98 49.71
Table 3: Performance metrics for test systems across different perturbations. SMAPE and MASE scores are reported as mean $\pm$ standard deviation.
Research Question Causal Diagram Metrics Used Comparison across Systems Comparison across Perturbations Key Conclusions
RQ1: Does $Z$ affect $R^{max}_{t}$, even though $Z$ has no effect on $P$? [Uncaptioned image] WRS {$S_{a}$: 3.96, $S_{g}$: 5.05, $S_{g}^{ni}$: 5.21, $S_{r}$: 5.34, $S_{p}^{ni}$: 5.38, $S_{c}$: 5.46, $S_{m}$: 5.75, $S_{p}$: 6.28, $S_{b}$: 6.9} {P2: 5.18, P1: 5.2, P3: 5.35, P0: 5.46} S with low statistical bias: $S_{a}$. S with high statistical bias: $S_{p}$. P that led to more statistical bias: P0 Analysis with more discrepancy: Inter-industry
RQ2: Does $Z$ affect the relationship between $P$ and $R^{max}_{t}$ when $Z$ has an effect on $P$? [Uncaptioned image] PIE % {$S_{g}^{ni}$: 523.09, $S_{g}$: 621.68, $S_{c}$: 742.19, $S_{a}$: 1206.44, $S_{m}$: 1466.13, $S_{p}^{ni}$: 1655.94, $S_{p}$: 2677.62, $S_{r}$: 2733.7, $S_{b}$: 5197.51} {P1: 995.19, P2: 1252.2, P3: 1563.94} S with low confounding bias: $S_{g}^{ni}$. S with high confounding bias: $S_{p}$. P that led to more confounding bias: P3. Confounder that led to more bias: Industry
RQ3: Does $P$ affect $R^{max}_{t}$ when $Z$ may have an effect on $R^{max}_{t}$? [Uncaptioned image] APE {$S_{g}^{ni}$: 6.49, $S_{g}$: 8.34, $S_{c}$: 8.42, $S_{m}$: 17.46, $S_{p}^{ni}$: 21.94, $S_{a}$: 28.28, $S_{p}$: 28.53, $S_{r}$: 29.65, $S_{b}$: 50.37} {P2: 12.74, P3: 17.34, P1: 21.11} S with low APE: $S_{g}^{ni}$. S with high APE: $S_{p}$. P with low APE: P2. P with high APE: P1. Confounder that led to high APE: Company
RQ4: Does $P$ affect the accuracy of $S$? This hypothesis does not necessitate a causal model for its evaluation. SMAPE, MASE, Sign Accuracy SMAPE: {$S_{c}$: 0.05, $S_{a}$: 0.06, $S_{g}$: 0.06, $S_{g}^{ni}$: 0.07, $S_{p}^{ni}$: 0.09, $S_{p}$: 0.1, $S_{m}$: 0.1, $S_{r}$: 0.83, $S_{b}$: 1.28}; MASE: {$S_{c}$: 4.67, $S_{g}$: 5.04, $S_{g}^{ni}$: 6.75, $S_{p}^{ni}$: 7.91, $S_{a}$: 8.54, $S_{p}$: 9.03, $S_{m}$: 9.13, $S_{r}$: 86.74, $S_{b}$: 947.56}; Sign Accuracy: {$S_{m}$: 40.91, $S_{p}$: 44.42, $S_{p}^{ni}$: 47.4, $S_{r}$: 49.71, $S_{g}$: 50.93, $S_{c}$: 51.99, $S_{g}^{ni}$: 52.94, $S_{a}$: 58.57, $S_{b}$: 62.6} SMAPE: {P3: 0.06, P0: 0.07, P1: 0.08, P2: 0.08}; MASE: {P3: 6.11, P0: 6.87, P2: 7.22, P1: 8.99}; Sign Accuracy: {P2: 49.26, P1: 49.42, P3: 49.73, P0: 49.97} S with good performance: $S_{c}$. S with poor performance: $S_{m}$. P with high impact on performance: P2.
Table 4: Summary of the research questions answered in the paper, causal diagram, metrics used in the experiment, average of the metric values compared across different systems, average computed across different perturbations, and the key conclusions drawn from the experiment. Overall, multi-modal FMTS demonstrated greater robustness and forecasting accuracy compared to multi-modal FMTS. TS FMTS demonstrated greater robustness and forecasting accuracy compared to GP FMTS. All the raw scores and ratings are shown in Table 2.
(a) SMAPE
(b) MASE
(c) Sign Accuracy
Figure 16: Bar plots showing the robustness metrics values across different systems and perturbations.
(a) WRS
(b) PIE % scores
(c) APE scores
Figure 17: Bar plots showing the robustness metrics values across different systems and perturbations.
Figure 18: Radar plots showing robustness metrics for all FMTS and $S_{a}$ under different perturbations.
Figure 19: Radar plots showing forecasting accuracy metrics for all systems under different perturbations.
Figure 20: Heatmap for all metrics for all models. Lighter shade indicates better performance.
Appendix F Additional User Study Results
In this section, we present all the hypotheses, the results from the statistical tests conducted to validate these hypotheses, and the conclusions drawn from the results.
Metric Q1 Q2 Q4 Q5 Q6 Q8 Q9 Q10 Q12 Q13 Q14
$\mu$ 3.1923 2.8077 2.5385 2.7692 2.9231 2.6923 2.9231 3.2308 2.6538 2.8077 3.0769
$\sigma$ 1.2335 1.3570 1.3336 1.1767 1.3834 1.0870 1.2625 1.4507 1.1981 1.3570 1.4676
t-statistic 4.9287 3.0349 2.0588 3.3333 3.4023 3.2476 3.7282 4.3259 2.7828 3.0349 3.7417
p-value 0.0000 0.0028 0.0250 0.0013 0.0011 0.0017 0.0005 0.0001 0.0051 0.0028 0.0005
Table 5: Summary of one sample right-tailed t-test results: Comparison of sample means to the hypothesized mean of 2 with a sample size of 26. The right-tailed p-values indicate whether the sample means are significantly greater the hypothesized mean. denotes that mean of responses for all the questions is greater than 2.
Hypothesis Test Performed Statistics Conclusion
There is a high positive correlation between users’ fairness rankings and rankings generated by our rating method. Spearman Rank Correlation $\rho=0.73$ The fairness rankings generated by our rating method aligns well with users’ rankings.
The mean of the responses for Q4 is less than or equal to the mean of the responses for Q6. Paired t-test t-statistic: -1.18, p-val: 0.12 Users found it easy to interpret the behavior of the systems from rankings compared to graphs and statistics with a confidence interval of 85 %.
There is a very high positive correlation between users’ rankings and rankings generated by our rating method. Spearman Rank Correlation $\rho$: 0.91 The robustness rankings generated by our rating method aligns very well with users’ rankings.
The mean of the responses for Q8 is less than or equal to the mean of the responses for Q10. Paired t-test t-statistic: -1.89, p-val: 0.03 Users found it easy to interpret the behavior of the systems from rankings compared to graphs and statistics with a confidence interval of 95 %.
There is a weak positive correlation between users’ rankings and rankings generated by our rating method. Spearman Rank Correlation $\rho$: 0.14 The robustness rankings generated by our rating method weakly aligns with users’ rankings.
The mean of the responses for Q12 is less than or equal to the mean of the responses for Q14. Paired t-test t-statistic: -1.62, p-val: 0.06 Users found it easy to interpret the behavior of the systems from rankings compared to graphs and statistics with a confidence interval of 90 %.
Table 6: Table with the hypotheses evaluated in the user study, statistical tests used to validate the hypotheses, results obtained, and conclusions drawn.
Appendix G Source Code for Data Processing
1# Convert data from Yahoo! finance to sliding window format.
2def sliding_window(data, window_size, company):
3 sequences = []
4 for i in range(len(data) - window_size):
5 seq = data[i:(i + window_size + 1)].tolist()
6 sequences.append([company] + seq)
7 return pd.DataFrame(sequences)
8
9# Perturbations:
10# Drop-to-zero: Every 80th stock price in the numerical data will be turned into zero.
11def drop_to_zero(df, col):
12
13 new_df = df.copy()
14 new_df.loc[new_df.index % 80 == 0, col] = 0
15
16 return new_df
17
18# Value halved: Every 80th stock price in the numerical data will be halved.
19def value_halved(df, col):
20
21 new_df = df.copy()
22 new_df.loc[new_df.index % 80 == 0, col] /= 2
23
24 return new_df
25
26# Missing values: Every 80th stock price in the numerical data will be NaN’.
27def missing_values(df, col):
28
29 new_df = df.copy()
30 new_df.loc[new_df.index % 80 == 0, col] = float(’nan’)
31
32 return new_df
33
34# Code to generate time series line plots.
35def plot_ts(input_path, output_path):
36 data = pd.read_csv(input_path)
37
38 companies = data.iloc[:, 0]
39 time_steps = data.iloc[:, 1:]
40
41 for i, company in enumerate(companies):
42 plt.figure(figsize=(12, 6))
43 plt.plot(time_steps.columns, time_steps.iloc[i], marker=’o’)
44 plt.title(fTime Series for {company}’, fontsize=19)
45 plt.xlabel(’Time Steps’, fontsize=17)
46 plt.ylabel(’Values’, fontsize=17)
47 plt.grid(True)
48 x_ticks = time_steps.columns[::5]
49 plt.xticks(x_ticks, rotation=45, fontsize=15)
50 plt.yticks(fontsize=15)
51 plt.tight_layout()
[…]
53 plt.close()
Appendix H Additional Implementation Details
All Forecasting Model Training Systems (FMTS) were executed on Colab notebooks utilizing the L4 GPU available through Colab Pro, which offers 22.5 GB of GPU RAM. Additional details regarding the models such as the inference times and other architectural details can be found in Section 5.1.
Hyperparameters set
  • MOMENT: head_dropout: 0.1, weight_decay: 0, freeze_encoder: True, freeze_embedder: True, freeze_head: False
  • Phi-3: _attn_implementation=’eager’, max_new_tokens: 300, temperature: 0.0, do_sample: False
  • Gemini: Temperature: 0. Rest of the parameters were default.
Appendix I Reproducibility Checklist
This paper:
  1. 1.
    Includes a conceptual outline and/or pseudocode description of AI methods introduced
    Answer: Yes (Appendix A).
  2. 2.
    Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results
    Answer: Yes (Section 5.2)
  3. 3.
    Provides well marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper
    Answer: Yes (Sections 1, 2)
    Does this paper make theoretical contributions? (yes/no)
    Answer: No
  4. 4.
    Does this paper rely on one or more datasets?
    Answer: Yes (Description in Section 5.1
  5. 5.
    A motivation is given for why the experiments are conducted on the selected datasets
    Answer: Yes (Sections 1, 2)
  6. 6.
    All novel datasets introduced in this paper are included in a data appendix.
    Answer: NA (We used an existing dataset from Yahoo! Finance.)
  7. 7.
    All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes.
    Answer: NA
  8. 8.
    All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations.
    Answer: NA
  9. 9.
    All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available.
    Answer: Yes (Section 5.1)
  10. 10.
    All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisfying.
    Answer: NA
  11. 11.
    Does this paper include computational experiments?
    Answer: Yes
  12. 12.
    Any code required for pre-processing data is included in the appendix.
    Answer: Yes, code required to convert the data downloaded from Yahoo! finance to sliding window, apply perturbations, and generate time series plots are provided in Appendix G.
  13. 13.
    All source code required for conducting and analyzing the experiments is included in a code appendix.
    Answer: Yes. You can find the source code and data here: https://anonymous.4open.science/r/rating-fmts-1B30/README.md
  14. 14.
    All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes.
    Answer: Yes
  15. 15.
    All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from
    Answer: Yes
  16. 16.
    If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results.
    Answer: Yes (provided in the source code).
  17. 17.
    This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks.
    Answer: Yes (in Appendix H)
  18. 18.
    This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics.
    Answer: Yes (Section 4.1 and Appendix C)
  19. 19.
    This paper states the number of algorithm runs used to compute each reported result.
    Answer: No.
  20. 20.
    Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information.
    Answer: Yes (Section 5.2, Appendix E, Section 5.4, and Appendix F)
  21. 21.
    The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank).
    Answer: Yes (Section 5.2, Appendix E, Section 5.4, and Appendix F)
  22. 22.
    This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes/partial/no/NA)
    Answer: Yes (Appendix Appendix H).
  23. 23.
    This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes/partial/no/NA)
    Answer: NA